OpenStudy (anonymous):

When finding the sum of the arithmetic sequence 4 + 1 + –2 + –5 + –8 + –11, you will have a step of work that contains the number –3. Use complete sentences to describe how the –3 was calculated and what it means.

OpenStudy (anonymous):

-3 is the common difference

OpenStudy (cwrw238):

1-4 = -3 -2-1 = -3 etc